For every integer n≥1,(n3+11n)⋅(8n−14n+27) is divisible by

6


7


42


84
1
Expert's answer
2020-11-02T20:56:52-0500

Hint : Since you probably missed some degrees in the expression "(8n-14n+27)" , I can only complete part of the task.


Using mathematical induction, we prove that "(n^3+11n)" is divisible by "6" for "n\in\mathbb{Z}^+" .

1 STEP: Induction basis

"n=1\to 1^3+11\cdot1=12=2\cdot6\,\,\,\vdots\,\,\,6"

2 STEP : Inductive guess

"\text{For all}\,\,\,k\le n\,\,\,(k^3+11k)\,\,\,\vdots\,\,\,6"



3 STEP : Inductive transition, it is necessary to prove that

"\text{For}\,\,\,k=n+1\to \left((n+1)^3+11(n+1)\right)\,\,\,\vdots\,\,\,6"

"(n+1)^3+11(n+1)=n^3+3n^2+3n+1+11n+11=\\[0.3cm]
=(n^3+11n)+(3n^2+3n)+12=\\[0.3cm]
=\underbrace{(n^3+11n)}_{\text{divided into 6 hypothesis}}+3n(n+1)+6\cdot 2"



It remains to understand why "3n(n+1)\,\,\,\vdots\,\,\,6" .

1 case : "n=2k"


"3n(n+1)=3\cdot(2k)\cdot(2k+1)=6\cdot\left(k(2k+1)\right)\,\,\,\vdots\,\,\,6"



2 case : "n=2k+1"


"3n(n+1)=3\cdot(2k+1)\cdot(2k+1+1)=6\cdot\left((k+1)(2k+1)\right)\,\,\vdots\,\,6"

Conclusion,


"(n+1)^3+11(n+1)=(n^3+11n)+3n(n+1)+6\cdot2\,\,\vdots\,\,6"

Q.E.D.